# Dirac quantum criticality in twisted double bilayer transition metal dichalcogenides

### v1

@@ IX Conclusions
We have investigated the ground-state phase diagram of $\boldsymbol{\Gamma}$-valley double bilayer TMDs at hole filling $\nu=2$ per moiré unit cell as a function of twist angle and applied pressure. At large twist angles, the system realizes a fully symmetric Dirac semimetal with spin-degenerate Dirac cones located at the moiré Brillouin-zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’. Reducing the twist angle drives a continuous transition into a Néel-type antiferromagnetic insulator on an emergent honeycomb lattice formed by the AB and BA stacking regions of the inner TMD layers. Unlike conventional antiferromagnets, where spin-density modulations occur on the microscopic lattice scale, the ordering wavelength of this antiferromagnetic state is set by the moiré length. At yet smaller twist angles, a level crossing occurs from the antiferromagnetic insulator to a ferromagnetic insulator with spin-split bands, stabilized by antichiral terms in the Hamiltonian.
@@
These theoretical predictions can be directly tested in hole-doped twisted double bilayer TMDs with a valence band edge at the $\boldsymbol{\Gamma}$ point, such as twisted double bilayer WSe22 [49]. In particular, our results suggest that the insulating state observed experimentally in twisted double bilayer WSe2 [35] is a Néel-type antiferromagnet with a spin-density wavelength set by the moiré scale. This substantial enhancement of the wavelength compared to conventional Néel antiferromagnets may enable direct imaging of the spin-density modulation using advanced nanoscale magnetometry techniques, including spin-resolved scanning tunneling microscopy [95], nitrogen-vacancy-center magnetometry [96], and potentially nanoSQUID devices [97, 98] or advanced magnetic force microscopy [99]. Pressure experiments on samples with twist angles just above the critical angle at ambient conditions could further test our predictions for the power-law scaling of the order parameter, $m_{\text{AFM}}\propto(p-p_{\mathrm{c}})^{\beta}$, with an unusually large exponent $\beta\approx 1.21$ compared to the conventional (2+1)D Heisenberg universality class. Transport experiments could likewise probe the predicted gap scaling, $\Delta\propto(p-p_{\mathrm{c}})^{\nu z}$, with $\nu\approx 1.20$ and $z=1$. Such measurements would establish the first direct experimental realization of a relativistic fermionic universality class.
@@
During the preparation of this manuscript, we became aware of a simultaneous study that employs a Hubbard-type model on the emergent honeycomb lattice as an effective description of twisted double bilayer WSe2 [100].
@@ Acknowledgements.
We thank Bilal Hawashin, Julian Kleeschulte, David Kurz, and Michael Scherer for valuable discussions. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 247310070 (SFB 1143, A07), Project No. 390858490 (Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147), and Project No. 411750675 (Emmy Noether program, JA2306/4-1).
@@
$|,\displaystyle\ket{\psi_{\mathbf{k},n}(\mathbf{r})}=\sum_{\begin{subarray}{c}\mathbf{G}\\ |\mathbf{G}|\leq 4|\mathbf{G}_{1,2}|\end{subarray}}\sum_{\ell}u_{\mathbf{k},n;\mathbf{G},\ell}\mathrm{e}^{\mathrm{i}(\mathbf{k}+\mathbf{G})\cdot\mathbf{r}}\ket{\ell}$, (A4)
@@ B.2 Strong-coupling Hamiltonian
$\mathcal{H}$ $=\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}\tilde{h}(\mathbf{k})c_{\mathbf{k}}-\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}:\mathrel{{\rho_{\mathbf{q}}\rho_{-\mathbf{q}}}}:-h_{\mathrm{HF}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k})$ (B10)
@@
$-\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}:\mathrel{\rho_{\mathbf{q}}\rho_{-\mathbf{q}}}:\,=\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}\rho_{\mathbf{q}}\rho_{-\mathbf{q}}+h_{\mathrm{F}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k}),$ (B11)
@@
$h_{\mathrm{H}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k})=\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}\left(\bar{\rho}_{\mathbf{q}}\rho_{-\mathbf{q}}+\bar{\rho}_{-\mathbf{q}}\rho_{\mathbf{q}}\right).$ (B12)
@@
B.3 Hartree-Fock treatment
The ground states we consider in the strong coupling limit are single Slater determinant states. We may therefore evaluate their energies using Hartree-Fock theory. For a given reduced density matrix $P$, this amounts to calculating the Hartree-Fock energy
@@ B.4 Energy functional
APPENDIX C CRITICAL PRESSURE FOR LARGER TWIST ANGLE
@@
In this appendix, we present additional data for the pressure-induced transition for a sample with a slightly larger twist angle than those in Fig. 5. Figure 8 shows the staggered magnetization density $m_{\text{AFM}}$ and the spectral gap $\Delta$, extrapolated to the thermodynamic limit, as functions of pressure $p$ for $\theta=2.8^{\circ}$. The results highlight the sensitivity of the critical pressure to the twist angle: For $\theta-\theta_{\mathrm{c}}(0)=0.05^{\circ}$, the semimetal-to-insulator transition occurs at $p_{\mathrm{c}}\simeq 0.2\,\text{GPa}$ (cf. Fig. 5), whereas for $\theta-\theta_{\mathrm{c}}(0)=0.10^{\circ}$, it shifts to $p_{\mathrm{c}}\simeq 0.6\,\text{GPa}$.

### v2

@@ IX Conclusions
We have investigated the ground-state phase diagram of $\boldsymbol{\Gamma}$-valley double bilayer TMDs at hole filling $\nu=2$ per moiré unit cell as a function of twist angle and applied pressure. At large twist angles, the system realizes a fully symmetric Dirac semimetal with spin-degenerate Dirac cones located at the moiré-Brillouin-zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’. Reducing the twist angle drives a continuous transition into a Néel-type antiferromagnetic insulator on an emergent honeycomb lattice formed by the AB and BA stacking regions of the inner TMD layers. Unlike conventional antiferromagnets, where spin-density modulations occur on the microscopic lattice scale, the ordering wavelength of this antiferromagnetic state is set by the moiré scale. At yet smaller twist angles, a level crossing occurs from the antiferromagnetic insulator to a ferromagnetic insulator with spin-split bands, stabilized by antichiral terms in the Hamiltonian.
@@
These theoretical predictions can be directly tested in hole-doped twisted double bilayer TMDs with a valence band edge at the $\boldsymbol{\Gamma}$ point. In particular, our results suggest that the insulating state observed experimentally in twisted double bilayer WSe2 [35] is a Néel-type antiferromagnet with a spin-density wavelength set by the moiré scale. This substantial enhancement of the wavelength compared to conventional Néel antiferromagnets may enable direct imaging of the spin-density modulation using advanced nanoscale magnetometry techniques, including spin-resolved scanning tunneling microscopy [101], nitrogen-vacancy-center magnetometry [102], and potentially nanoSQUID devices [103, 104] or advanced magnetic force microscopy [105]. Pressure experiments on samples with twist angles just above the ambient-condition critical angle could further test our predictions for the power-law scaling of the order parameter, $m_{\text{AFM}}\propto(p-p_{\mathrm{c}})^{\beta}$, with an unusually large exponent $\beta\approx 1.21$ compared to the conventional (2+1)D Heisenberg universality class. Transport experiments could likewise probe the predicted gap scaling, $\Delta\propto(p-p_{\mathrm{c}})^{\nu z}$, with $\nu\approx 1.20$ and $z=1$. Ultrafast optical techniques [106, 107] may enable tests of the predicted crossover from Gross-Neveu-Heisenberg to Heisenberg universality, reflected in the frequency scaling of the dynamical spin structure at low temperatures in weakly strained samples.
From a broader perspective, two additional remarks are in order. First, quantum critical points involving gapless fermionic excitations have been discussed previously in a wide range of strongly-correlated systems, including cuprate and iron-based superconductors, heavy-fermion compounds, and other moiré materials. In these settings, however, the absence of a sufficiently high symmetry has precluded a controlled theoretical description of the universal physics governing the transition. By contrast, the emergent Lorentz symmetry that we predict at the semimetal-to-insulator transition in twisted double bilayer WSe2 enables a field-theoretical treatment that is fully accessible within a standard $\varepsilon$ expansion about the upper critical spatial dimension $d=3$. The quantum critical point in twisted double bilayer WSe2 may therefore constitute the first experimentally realizable example of a fermionic quantum universality class that is under complete theoretical control. Second, a variant of the Gross-Neveu-Heisenberg model identified here, incorporating randomized couplings, has recently been shown to support a superconducting dome emerging from strong critical fluctuations as a secondary instability above the quantum critical point [108, 109]. While we do not expect such a secondary instability to arise in pristine twisted double bilayer WSe2, an important direction for future work is to identify the conditions under which randomized Gross-Neveu-Heisenberg criticality may be realized, potentially offering new insights into the emergence of strong-coupling superconductivity above quantum critical points.
@@
After this manuscript was submitted, a simultaneous study appeared on arXiv, which employs a Hubbard-type model on the emergent honeycomb lattice as an effective description of twisted double bilayer WSe2 [110].
@@ Acknowledgements.
We thank Bilal Hawashin, Julian Kleeschulte, David Kurz, and Michael Scherer for valuable discussions. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 247310070 (SFB 1143, A07), Project No. 390858490 (Würzburg-Dresden Cluster of Excellence ctd.qmat, EXC 2147), and Project No. 411750675 (Emmy Noether program, JA2306/4-1).
@@
$|,\displaystyle\ket{\psi_{\mathbf{k},n}(\mathbf{r})}=\frac{1}{\sqrt{(2\pi)^{2}}}\sum_{\begin{subarray}{c}\mathbf{G}\\ |\mathbf{G}|\leq 4|\mathbf{G}_{1,2}|\end{subarray}}\sum_{\ell}u_{\mathbf{k},n;\mathbf{G},\ell}\mathrm{e}^{\mathrm{i}(\mathbf{k}+\mathbf{G})\cdot\mathbf{r}}\ket{\ell}$, (A4)
@@ B.2 Strong-coupling Hamiltonian
$\mathcal{H}$ $=\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}\tilde{h}(\mathbf{k})c_{\mathbf{k}}-\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}:\mathrel{{\rho_{\mathbf{q}}\rho_{-\mathbf{q}}}}:-\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}h_{\mathrm{HF}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k})c_{\mathbf{k}}$ (B10)
@@
$-\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}:\mathrel{\rho_{\mathbf{q}}\rho_{-\mathbf{q}}}:\,=\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}\rho_{\mathbf{q}}\rho_{-\mathbf{q}}+\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}h_{\mathrm{F}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k})c_{\mathbf{k}},$ (B11)
@@
$\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}h_{\mathrm{H}}[\tfrac{1}{2}\mathbb{1}](\mathbf{k})c_{\mathbf{k}}=\frac{1}{2A}\sum_{\mathbf{q}}V_{\mathbf{q}}\left(\bar{\rho}_{\mathbf{q}}\rho_{-\mathbf{q}}+\bar{\rho}_{-\mathbf{q}}\rho_{\mathbf{q}}\right).$ (B12)
@@
B.3 Energy scales
To confirm the validity of our strong-coupling analysis, we numerically estimate the relevant energy scales. To this end, we explicitly evaluate the chiral part $\Lambda_{\mathrm{S}}(\mathbf{k},\mathbf{q})$ and the antichiral part $\Lambda_{\mathrm{A}}(\mathbf{k},\mathbf{q})$ of the overlap matrices, their magnitudes $F_{\mathrm{S}}(\mathbf{k},\mathbf{q})$ and $F_{\mathrm{A}}(\mathbf{k},\mathbf{q})$, and the corresponding interaction energy scales $U_{S}$ and $U_{A}$. The latter are then compared with the average bandwidth $\bar{t}$, which represents the effective kinetic energy scale.
Note that the definition of the quantities $\Lambda_{\mathrm{S}/\mathrm{A}}(\mathbf{k},\mathbf{q})$ involves the sublattice operator $\sigma_{z}$, but there exists no explicit expression for the matrix elements of this operator in the band basis, because the sublattice degree of freedom is an emergent property. In our numerics, we circumvent this problem by extracting the sublattice operator $\sigma_{z}$ from the density matrix $P_{\text{AFM}}$ of an antiferromagnetic state: In the absence of dispersion, converging the antiferromagnetic state yields a density matrix of the form $P_{\text{AFM}}=\frac{1}{2}(\mathbb{1}+\sigma_{z}\mathbf{n}\cdot\mathbf{s})$, where we can fix $\mathbf{n}=\mathbf{e}_{z}$ by choosing a suitable initialization and then solve for $\sigma_{z}$.
We extract the magnitudes $F_{\mathrm{S}/\mathrm{A}}(\mathbf{k},\mathbf{q})$ of the overlap matrices from the basis-independent expression
[…]
The corresponding energy scales $U_{\mathrm{S}/\mathrm{A}}$ of the interaction terms $\mathcal{H}_{\mathrm{S}/\mathrm{A}}$ may then be estimated by [54]
$U_{\mathrm{S}}$ $=\frac{1}{2AL^{2}}\sum_{\mathbf{k},\mathbf{q}}V_{\mathbf{q}}F_{\mathrm{S}}(\mathbf{k},\mathbf{q})^{2},$ (B15)
$U_{\mathrm{A}}$ $=\frac{1}{2AL^{2}}\sum_{\mathbf{k},\mathbf{q}}V_{\mathbf{q}}\left(2F_{\mathrm{S}}(\mathbf{k},\mathbf{q})F_{\mathrm{A}}(\mathbf{k},\mathbf{q})+F_{\mathrm{A}}(\mathbf{k},\mathbf{q})^{2}\right).$ (B16)
These quantities are compared to the $\mathbf{k}$-averaged bandwidth $\bar{t}$ of the renormalized dispersion $\tilde{h}(\mathbf{k})$, given by
$\bar{t}=\frac{1}{2L^{2}}\sum_{\mathbf{k}}\left(\varepsilon_{+}(\mathbf{k})-\varepsilon_{-}(\mathbf{k})\right),$ (B17)
where $\varepsilon_{\pm}(\mathbf{k})$ is the upper (lower) band in the spectrum of $\tilde{h}(\mathbf{k})$.
Figure 9 displays the $\mathbf{k}$-averaged bandwidth $\bar{t}$ and the antichiral interaction energy scale $U_{\mathrm{A}}$, both normalized to the chiral interaction energy scale $U_{\mathrm{S}}$, together with the ratio of the average magnitudes $\overline{F_{\mathrm{A}}}/\overline{F_{\mathrm{S}}}$ of the antichiral and chiral components of the overlap matrices. Here, the averages are defined as $\overline{F_{\mathrm{S}/A}}=\sum_{\mathbf{k},\mathbf{q}}F_{\mathrm{S}/A}(\mathbf{k},\mathbf{q})/L^{4}$. These results confirm that for small twist angles, both the dispersion and the antichiral interaction can indeed be treated as small perturbations to the chiral Hamiltonian $\mathcal{H}_{\mathrm{S}}$.
Figure 9: Kinetic energy scale $\bar{t}$ (blue dots) and antichiral interaction energy scale $U_{\mathrm{A}}$ (red crosses), both normalized to the chiral interaction energy scale $U_{\mathrm{S}}$, together with the ratio of the average magnitudes $\overline{F_{\mathrm{A}}}/\overline{F_{\mathrm{S}}}$ of the antichiral and chiral components of the overlap matrices (orange triangles). Calculations were performed using a fixed effective permittivity $\epsilon_{\text{eff}}=110$, on an $18\times 18$ momentum-space grid, keeping two bands per spin species. At small twist angles, the chiral interaction dominates over both the antichiral contribution and the dispersion, justifying a perturbative treatment. Near the critical twist angle $\theta\simeq 2.7^{\circ}$, the kinetic energy scale $\overline{t}$ and the interaction energy scale $U_{\mathrm{S}}$ are comparable, as expected.
B.4 Hartree-Fock treatment
The ground states we consider in the strong-coupling limit are single Slater determinant states. We may therefore evaluate their energies using Hartree-Fock theory. For a given reduced density matrix $P$, this amounts to calculating the Hartree-Fock energy
@@ B.5 Energy functional
B.6 Effect of dispersion
In the following, we perturbatively analyze the effect of the renormalized kinetic energy term
$\mathcal{H}_{0}=\sum_{\mathbf{k}}c^{\dagger}_{\mathbf{k}}\tilde{h}(\mathbf{k})c_{\mathbf{k}}$ (B29)
in the two-band model [Eq. (18)] on the energy competition between the magnetic orderings. The first-order energy correction turns out to be identical for both phases, but the second-order contribution favors the antiferromagnetic ground state. Intuitively, this can be understood as an analogue of the antiferromagnetic Heisenberg exchange interaction with coupling $J\propto t^{2}/U$ that arises in the strong-coupling expansion of the Hubbard model.
Note that the $C_{2z}\mathcal{T}$ and SU(2) symmetries imply the form
$\tilde{h}(\mathbf{k})=(a(\mathbf{k})\sigma_{0}+b(\mathbf{k})\sigma_{x}\mathrm{e}^{\mathrm{i}\Phi_{0}(\mathbf{k})\sigma_{z}})s_{0}$ (B30)
for the matrix $\tilde{h}(\mathbf{k})$, where $a$, $b$ and $\Phi_{0}$ are real-valued functions of the wavevector $\mathbf{k}$. The four-dimensional ground-state subspace $\mathcal{D}$ of the chiral Hamiltonian $\mathcal{H}_{\mathrm{S}}$ is spanned by the states
[…]
which diagonalize the perturbation $\mathcal{H}_{0}$ on $\mathcal{D}$ and thus represent a suitable basis for the degenerate perturbation theory. The corresponding energy corrections to first order in $\mathcal{H}_{0}$ are given by
[…]
For the antiferromagnetic states, we instead find
[…]
with electron-hole pair states given by
[…]
and
[…]
APPENDIX C EFFECTIVE PERMITTIVITY
This appendix provides a theoretical estimate for the effective permittivity $\epsilon_{\mathrm{eff}}(2.7^{\circ})$ at the critical twist angle, showing that values $\epsilon_{\text{eff}}\sim\mathcal{O}(100)$ are indeed realistic for twisted double bilayer WSe2. We first evaluate the momentum-dependent dielectric function $\epsilon(\mathbf{q})$ within the random phase approximation, and subsequently calculate an effective constant permittivity $\epsilon_{\text{eff}}$ in terms of a suitably weighted average of $\epsilon(\mathbf{q})$.
Following Ref. [59], we assume that the diagonal elements of the dielectric tensor dominate and can be evaluated within the random phase approximation as
$\epsilon(\mathbf{q})=\epsilon_{\mathrm{env}}+v(\mathbf{q})\Pi_{0}(\mathbf{q}).$ (B43)
Here, $\epsilon_{\mathrm{env}}$ accounts for the substrate contribution, $v(\mathbf{q})=e^{2}/2\epsilon_{0}|\mathbf{q}|$ is the bare Coulomb potential, and the independent-particle polarizability $\Pi_{0}$ is given by [59]
[…]
To facilitate the theoretical modeling, we further decompose the dielectric function as
$\epsilon(\mathbf{q})=\epsilon_{\mathrm{env}}+v(\mathbf{q})(\Pi_{0}^{\mathrm{cRPA}}(\mathbf{q})+\Pi_{0}^{\mathrm{flat}}(\mathbf{q})),$ (B46)
where $\Pi_{0}^{\mathrm{cRPA}}(\mathbf{q})$ accounts for transitions involving remote bands, and
[…]
arises from transitions within the two nearly-flat graphene-like bands $\varepsilon_{\pm,\mathbf{q}}$. From previous calculations in the context of twisted bilayer graphene [77, 59, 31], we expect the flat-band contribution $\Pi_{0}^{\text{flat}}$ to yield the dominant contribution to the effective permittivity. It can be directly computed within the low-energy continuum approximation. Accurately computing the remote-band contribution $\Pi_{0}^{\text{cRPA}}$ requires to go beyond the continuum approximation, as also remote bands that are not well captured within the continuum model may contribute to the effective permittivity. Here, we estimate this contribution from the permittivity of an uncoupled WSe2 tetralayer, $v(\mathbf{q})\Pi_{0}^{\text{cRPA}}(\mathbf{q})\simeq g_{\mathrm{l}}(\epsilon_{\text{mono-WSe${}_{2}$}}-1)$, where $\epsilon_{\text{mono-WSe${}_{2}$}}\simeq 2.9$ is the in-plane static permittivity of a WSe2 monolayer [114], $g_{\mathrm{l}}=4$ counts the number of layers, and we subtract one to prevent double counting of the vacuum permittivity. This assumption neglects the increase of screening that arises from the flattening of the remote bands for small twist angles and may thus be understood as a lower bound for the corresponding contribution from $\Pi_{0}^{\text{cRPA}}$ to $\epsilon(\mathbf{q})$ [77, 59]. In passing, we note that a meaningful comparison of the critical interaction strength in twisted double bilayer TMDs with the critical value $(U_{0}/t_{1})_{\mathrm{c}}\simeq 3.8$ for the honeycomb-lattice Hubbard model [4] requires accounting for the screening captured by $\Pi_{0}^{\text{cRPA}}$. Including this effect substantially lowers the critical Wigner-Seitz radius relative to the estimate given in Ref. [35].
Figure 10: Effective permittivity $\epsilon_{\mathrm{eff}}$ as a function of the twist angle $\theta$ in the Dirac semimetal phase for a substrate permittivity $\epsilon_{\mathrm{env}}=5$. Here, a $17\times 17$ momentum-space grid is used to avoid singularities in the momentum sum of Eq. (B47). At the experimentally determined critical twist angle $\theta_{\mathrm{c}}=2.7^{\circ}$, we find $\epsilon_{\mathrm{eff}}(\theta_{\mathrm{c}})\simeq 82$.
An effective dielectric constant suitable as an input for the Hartree-Fock analysis can be obtained by averaging the momentum-dependent dielectric function $\epsilon(\mathbf{q})$ such that the interaction energy scale $U$ is preserved. In analogy with Appendix B.3, we define
$U=\frac{1}{2AL^{2}}\sum_{\mathbf{q},\mathbf{k}}\tilde{V}_{\mathbf{q}}F(\mathbf{k},\mathbf{q})^{2}\,$, (B48)
[…]
$\epsilon_{\mathrm{eff}}=\frac{\sum_{\mathbf{q},\mathbf{k}}\epsilon(\mathbf{q})\tilde{V}_{\mathbf{q}}\,F(\mathbf{k},\mathbf{q})^{2}}{\sum_{\mathbf{q},\mathbf{k}}\tilde{V}_{\mathbf{q}}\,F(\mathbf{k},\mathbf{q})^{2}}\,.$ (B49)
Numerical results for the resulting effective dielectric constant are shown in Fig. 10. At the experimentally determined critical twist angle $\theta_{\mathrm{c}}=2.7^{\circ}$, we find $\epsilon_{\mathrm{eff}}(\theta_{\mathrm{c}})\simeq 82$, close to the semimetal-to-insulator phase boundary shown in Fig. 4(c). We note that the above calculation is based on the noninteracting continuum Hamiltonian assuming a Dirac semimetal ground state. Interactions may modify the effective permittivity. This effect is expected to be particularly significant in the presence of a finite interaction-induced band gap, which is why we only present results for twist angles above approximately $\theta_{\mathrm{c}}$.
APPENDIX D CRITICAL PRESSURE FOR LARGER TWIST ANGLE
@@ B.6 Effect of dispersion
This appendix presents additional data on the pressure-induced transition for a sample with a slightly larger twist angle than in Fig. 5. Figure 11 shows the staggered magnetization density $m_{\text{AFM}}$ and the spectral gap $\Delta$, extrapolated to the thermodynamic limit, as functions of pressure $p$ for $\theta=2.8^{\circ}$. The results highlight the sensitivity of the critical pressure to the twist angle: For $\theta-\theta_{\mathrm{c}}(0)=0.05^{\circ}$, the semimetal-to-insulator transition occurs at $p_{\mathrm{c}}\simeq 0.2\,\text{GPa}$ (cf. Fig. 5), whereas for $\theta-\theta_{\mathrm{c}}(0)=0.10^{\circ}$, it shifts to $p_{\mathrm{c}}\simeq 0.6\,\text{GPa}$.